# VisioPhysioENet: Multimodal Engagement Detection using Visual and Physiological Signals

### v1

@@ Abstract
This paper presents VisioPhysioENet, a novel multimodal system that leverages visual cues and physiological signals to detect learner engagement. It employs a two-level approach for visual feature extraction using the Dlib library for facial landmark extraction and the OpenCV library for further estimations. This is complemented by extracting physiological signals using the plane-orthogonal-to-skin method to assess cardiovascular activity. These features are integrated using advanced machine learning classifiers, enhancing the detection of various engagement levels. We rigorously evaluate VisioPhysioENet on the DAiSEE dataset, where it achieves an accuracy of 63.09%, demonstrating a superior ability to discern various levels of engagement compared to existing methodologies. The proposed system’s code can be accessed at github.com/MIntelligence-Group/VisioPhysioENet. footnotetext: denotes equal contribution, # denotes corresponding author.
@@ I Introduction
Engagement detection in educational settings is essential for optimizing learning outcomes [1]. Traditional methods, often limited by their reliance on a single data type, may not fully capture the complexities of learner engagement [2, 3]. Recent studies advocate for multimodal engagement detection, which integrates various data types to overcome the limitations of unimodal approaches [4, 5]. Although these multimodal systems combining visual and physiological signals show improved accuracy and comprehensiveness [6, 7], they still face challenges such as data imbalance and limited generalizability across different datasets [8, 9]. Additionally, most research has predominantly focused on visual data, with less emphasis on exploiting multimodality [2].
@@
The proposed multimodal engagement detection framework, VisioPhysioENet, addresses these challenges by leveraging advanced machine learning algorithms and multimodal data integration. This ensures a balance between accuracy and computational efficiency, avoiding the complexities of deep neural architectures [10, 9]. The system employs a two-level approach for visual feature extraction: first, using Dlib [11] to extract facial landmarks, and OpenCV [12] for eye and head metrics like Eye Aspect Ratio (EAR), Pitch, Yaw, and Roll. These are transformed into numerical features and categorized into actionable classes, including eye openness, gaze direction, and head position. In parallel, remote photoplethysmography (rPPG) signals are captured through video inputs to monitor cardiovascular activity. A key innovation of our work is using rPPG to extract physiological signals non-invasively without measurement devices, offering a convenient approach to monitoring cardiovascular activity.
@@
Transitioning from the initial processing of time-serialized signals to handling single-value data points in the final phase, VisioPhysioENet significantly reduces computational demands and enhances system robustness. This integration of visual and physiological data through advanced machine learning models, specifically multi-output classifiers, and the application of late fusion techniques ensures improved accuracy and robustness in detecting engagement levels. When rigorously evaluated on the Dataset for Affective States in E-Environments (DAiSEE) dataset, VisioPhysioENet achieves an accuracy of 63.09%, demonstrating a superior ability to discern various levels of engagement compared to existing methodologies. The major contributions of this paper are listed as follows:
@@
  • Proposal of VisioPhysioENet that combines physiological information with visual data without using additional measurement devices. It is lightweight, enabling fast processing with minimal computational load.
@@ II Related Works
The detection of learner engagement has significantly evolved, initially focusing on single data types. Abedi et al. [6] combined ResNet with temporal convolutional networks to enhance accuracy in datasets like DAiSEE. Selim et al. [7] used EfficientNet and Long Short Term Memory (LSTM) network for video data analysis but face data imbalance issue.
Physiological signal integration has been explored by Sarkar et al. [13] who used ECG data for emotion classification, and Gupta et al. [10] focused on facial cues for engagement detection. Studies by Xie et al. [8] and Mehta et al. [9] show advancements in multimodal data analysis, stressing the need for better integration techniques in real-world settings.
Contributions from Vedernikov et al. [14] and Dubbaka et al. [15] have refined architectures for diverse settings. Work by Copur et al. [16] and Singh et al. [17] expand engagement detection into informal and digital environments, while Sherly et al. [18], Hasnine et al. [19], and Sharma et al. [20] have integrated various machine learning models to enrich the field.
Chen et al. [3] and Lee et al. [21] addressed real-time challenges in engagement systems, focusing on robustness under varied conditions. Gupta et al. [22] provided real-time deep learning solutions. Efforts by Yang et al. [23] and Liao et al. [24], along with the work by Savchenko et al. [25] and Abedi et al. [26], demonstrate the necessity for systems like VisioPhysioENet to effectively synthesize visual and physiological data for nuanced behaviour analysis.
Building on these insights, our research addresses critical gaps by implementing a novel multimodal engagement detection system, VisioPhysioENet, which is lightweight and enables faster processing with minimal computational overhead. Our approach incorporates advanced multimodal fusion techniques, including early and late fusion strategies, significantly enhancing robustness and accuracy in engagement detection across diverse educational settings.
@@ III VisioPhysioENet
The architecture of the proposed system, VisioPhysioENet, is shown in Fig. 1 and described below.
Figure 1: Schematic architecture of VisioPhysioENet encompassing feature extraction modules for visual and physiological modalities along with the fusion module.
@@ III-A Visual Feature Extraction Phase
The visual feature extraction model operates at two levels, denoted as £$\pounds_{i}^{V}$, to extract features from videos ($V$) by processing frames ($f_{i}$) sequentially. The first level, £$\pounds_{1}^{V}$, involves extracting direct features as facial landmarks ($p_{1}$ to $p_{68}$) including eye metrics points ($\varepsilon$) and head position estimation points ($\hbar$). These features are identified using the 68-FLM landmarks detection model provided by the Dlib library.
@@
The £$\pounds_{1}^{V}$ also involves extracting numerical features from Eq (1)’s data. This includes the Eye Aspect Ratio ($\xi$) and parameters related to the position of the head $[HP]$, such as pitch $\theta_{x}$, yaw $\theta_{y}$, and roll $\theta_{z}$. The EAR ($\xi$) is calculated using the facial landmark points of both the left ($\xi_{L}$) and right eyes ($\xi_{R}$), and then their mean is taken.
@@
where $/2]{\sum}$ denotes the computation of the mean and assignment operator. For head pose ($HP$) estimation, specific facial landmarks, such as nose tip, chin, corners of the eye, and mouth corners are used. The process assumes no camera distortions during focal length estimation. The rotation vector $\mathbf{RV}_{[i,j]}^{3\times 1}\big{\}}$, describing the object’s orientation relative to the camera, is computed using the solvePNP function. The OpenCV first extracts the axis ‘u’ and angle $\theta$ from the rotation vector $\mathbf{RV}_{[i,j]}^{3\times 1}\big{\}}$. It then constructs the skew-symmetric matrix K and calculates the R using the below formula.
@@
The next level, i.e., £$\pounds_{2}^{V}$, involves extracting categorical variables such as eye gaze direction ($\varphi$), head position ($\hbar$), and eye openness category ($\epsilon$) from the numerical variables derived in £$\pounds_{1}^{V}$. To calculate $\varphi$, the image center($\odot$) and eye center ($\otimes$) have been used, as follows:
@@
The gaze direction ($\varphi$), computed based on the gaze vector consisting of horizontal (x) and vertical (y) components, indicates whether a person is looking left, right, up, down, or forward based on its values. $\hbar$, computed based on $\theta_{x},\theta_{y}$, and $\theta_{z}$, has 7 classes, including neutral, tilted, or turned in various directions. Eye openness ($\epsilon$), measured by EAR ($\xi$), is categorized into three classes: fully open, partially closed, and closed. All categorical variables depend on thresholds, which were determined experimentally for each variable.
@@ III-B Physiological Feature Extraction Phase
The second phase of feature extraction focuses on extracting the Blood Volume Pulse (BVP) signal from video frames to reflect cardiovascular activity using the Plane-Orthogonal-to-Skin (POS) method, a non-invasive technique for rPPG. This method capitalizes on subtle skin colour changes caused by blood flow to estimate the BVP signal.
The £$\pounds_{1}^{P}$ denotes the first level in extracting physiological features, which begins with extracting signals ($S$) by calculating the average RGB values from a specified region of interest (ROI) i.e. forehead $(\Re_{f})$ for the proposed model in each video frame.
@@
where $R_{i}(t),G_{i}(t),$ and $B_{i}(t)$ are the RGB values from the region of interest in the $i^{th}$ video frame, and N is the total number of pixels in the $(\Re_{f})$. Then, temporal normalization is applied to the RGB values by dividing each colour signal by its mean over a specific time window. This transformation emphasizes the pulsatile component, primarily captured in the green channel, which is most sensitive to blood flow changes.
@@
To effectively apply the POS algorithm, we first need to identify the Region of Interest (ROI) i.e. ($\Re_{f}$). Finally, to extract the BVP ($\mathbf{B}$) signal, the POS algorithm has been applied to the extracted $\Re_{f}$.
$\text{B}(t)=POS_{Algo}(\text{RGB}(t))$ (8)
@@
To calculate the $RR_{i}$ interval, analysis has been done to identify peaks representing heartbeats $(t_{i+1}-t_{i})$. From the mean value $\overline{RR}$, heart rate in beats per minute (BPM) is given by:
$\text{BPM}=\frac{\overline{RR}}{60}$ (10)
Some more features in £$\pounds_{2}^{P}$, like average Peak-to-Peak Interval $(PPI)$ to assess heart rhythm regularity, are calculated as follows:
$\text{PPI}_{\text{avg}}=\frac{1}{N}\sum_{i=1}^{N}RR_{i}$ (11)
Similarly, from $RR_{i}$, the cumulative sums of systolic $(CS_{Sys})$ and diastolic $(CS_{Dia})$ peaks are extracted. The average heart rate over the entire signal duration is computed, and overall heart rate trends are summarized based on the BVP signal.
[…]
where $\overline{RR}$ is the mean RR interval over the entire recording period. $\text{BPM}(t)$ is the instantaneous heart rate at time $t$, and $T$ is the total duration of the signal.

### v2

@@ Abstract
This paper presents VisioPhysioENet, a novel multimodal system that leverages visual and physiological signals to detect learner engagement. It employs a two-level approach for extracting both visual and physiological features. For visual feature extraction, Dlib is used to detect facial landmarks, while OpenCV provides additional estimations. The face_recognition library, built on Dlib, is used to identify the facial region of interest specifically for physiological signal extraction. Physiological signals are then extracted using the plane-orthogonal-to-skin method to assess cardiovascular activity. These features are integrated using advanced machine learning classifiers, enhancing the detection of various levels of engagement. We thoroughly tested VisioPhysioENet on the DAiSEE dataset. It achieved an accuracy of 63.09%. This shows it can better identify different levels of engagement compared to many existing methods. It performed 8.6% better than the only other model that uses both physiological and visual features.
@@ I Introduction
Engagement detection in educational settings is essential for optimizing learning outcomes [1]. Traditional methods, often limited by their reliance on a single data type, may not fully capture the complexities of learner engagement [2, 3]. Recent studies advocate for multimodal engagement detection, which integrates various data types to overcome the limitations of unimodal approaches [4, 5]. Although these multimodal systems combining visual and physiological signals show improved accuracy and comprehensiveness [6, 7], they still face challenges such as data imbalance and limited generalizability across different datasets [8, 9]. Additionally, most of the research in this domain has predominantly focused on visual data, with little emphasis on exploiting multimodality [2]. VisioPhysioENet utilizes a multimodal approach to uncover subtle patterns of disengagement, enabling precise interventions like adapting instructional methods. This strategy enhances learning and provides educators with tools for meaningful academic success.
@@
The proposed multimodal engagement detection framework, VisioPhysioENet, addresses these challenges by leveraging advanced machine learning algorithms and multimodal data integration. This ensures a balance between accuracy and computational efficiency, avoiding the complexities of deep neural architectures [10, 9]. The system employs a two-level approach for visual and physiological feature extraction. First, using Dlib [11] to extract facial landmarks, and OpenCV [12] for eye and head metrics like Eye Aspect Ratio (EAR), Pitch, Yaw, and Roll. These are transformed into numerical features and categorized into actionable classes, including eye openness, gaze direction, and head position. In parallel, remote photoplethysmography (rPPG) signals [13] are captured from video input to monitor cardiovascular activity. The rPPG method utilizes the face_recognition[14] library, which incorporates Dlib, to identify and extract the region of interest (ROI) for processing. These extracted signals are then further analyzed to compute meaningful metrics, such as heart rate, average peak-to-peak interval time, cumulative sum of systolic peaks, and cumulative sum of diastolic peaks.
@@
Transitioning from the initial processing of time-serialized signals to handling single-value data points in the final phase, VisioPhysioENet significantly reduces computational demands and enhances system robustness. This integration of visual and physiological data through advanced machine learning models, specifically multi-output classifiers, and the application of late fusion techniques ensures improved accuracy and robustness in detecting engagement levels. When rigorously evaluated on the Dataset for Affective States in E-Environments (DAiSEE) dataset [15], VisioPhysioENet achieves an accuracy of 63.09%, demonstrating a superior ability to discern various levels of engagement compared to existing methodologies. The major contributions of this paper are listed as follows:
@@
  • Proposal for VisioPhysioENet, a framework that integrates physiological information with visual data, eliminating the need for additional measurement devices while delivering a lightweight, fast, and computationally efficient solution.
@@ II Related Works
The detection and analysis of learner engagement have advanced significantly, progressing from single-modal approaches to sophisticated multimodal systems. These methods aim to tackle the complexities of engagement detection in dynamic, real-world scenarios.
Initial works focused on single data modalities, such as visual or physiological signals. Abedi et al. [6] employed ResNet with temporal convolutional networks to extract spatiotemporal features from video datasets like DAiSEE. This approach highlighted the value of leveraging temporal patterns for engagement detection. Similarly, Selim et al. [7] utilized EfficientNet combined with Long Short-Term Memory (LSTM) networks for video-based engagement detection, showcasing the effectiveness of sequential modelling. However, their efforts were constrained by imbalanced datasets, a recurring challenge in engagement research.
The integration of physiological data has added depth to engagement analysis. Sarkar et al. [16] demonstrated the utility of ECG data for emotion classification, which directly correlates with engagement levels. Gupta et al. [10] further extended this domain by combining facial features with physiological data such as heart rate variability, enabling a more holistic assessment of learner engagement. These efforts underline the importance of physiological signals in capturing subtle, non-visible aspects of engagement.
To overcome the limitations of single-modal approaches, researchers have increasingly focused on multimodal fusion techniques. Xie et al. [8] and Mehta et al. [9] explored early and late fusion strategies to integrate visual and physiological data streams. Early fusion combines raw features from multiple modalities, offering a unified feature representation, while late fusion integrates modality-specific predictions to enhance decision-making. These techniques have proven effective in capturing nuanced behavioral patterns and improving the robustness of engagement detection in complex environments.
Recent advancements in architecture design have further improved engagement detection. Vedernikov et al. [17] and Dubbaka et al. [18] refined deep learning models to handle diverse educational settings. Copur et al. [19] and Singh et al. [20] extended engagement analysis into informal and digital environments, emphasizing the adaptability of these systems.
Machine learning models continue to play a pivotal role in engagement analysis. Sherly et al. [21], Hasnine et al. [22], and Sharma et al. [23] implemented various machine learning techniques to address challenges such as data imbalance and computational efficiency. Ensemble methods, including random forests and boosting algorithms, have been particularly effective in improving prediction accuracy by combining multiple weak learners into a robust model.
Real-time engagement detection has been a critical focus area. Chen et al. [3] and Lee et al. [24] addressed the challenges of deploying engagement detection systems in real-time, emphasizing the importance of lightweight models that can operate efficiently under varied conditions. Gupta et al. [25] proposed deep learning-based systems tailored for dynamic environments, showcasing the potential of these models to adapt to changing engagement levels.
The synthesis of visual and physiological data has led to groundbreaking systems like VisioPhysioENet. Yang et al. [26] and Liao et al. [27] demonstrated the effectiveness of combining modalities for nuanced behavior analysis. Savchenko et al. [28] introduced streamlined CNN architectures optimized for engagement detection, while Abedi et al. [29] highlighted the capabilities of engagement analysis for real-time applications.
Building on these advancements, our research addresses critical gaps in engagement detection. We propose VisioPhysioENet, a novel multimodal system that incorporates advanced fusion techniques, including early and late fusion, to seamlessly integrate visual and physiological data. The system converts unstructured input data (video frames) into structured numeric tabular data, achieving dimensionality reduction where traditional machine learning models excel.[30] It is optimized for lightweight deployment, ensuring minimal computational overhead while maintaining high accuracy. By focusing on real-time processing and scalability, VisioPhysioENet provides robust engagement detection across diverse educational and informal learning environments.
@@ III VisioPhysioENet
The architecture of the proposed system, VisioPhysioENet, is shown in Fig. 1 and described below. The theory for feature extraction and fusion is presented as follows.
Figure 1: Schematic architecture of VisioPhysioENet encompassing feature extraction modules for visual and physiological modalities and the fusion module.
@@ III-A Visual Feature Extraction Phase
The visual feature extraction model operates at two levels, denoted as £$\pounds_{i}^{V}$, to extract features from videos ($V$) by processing frames ($f_{i}$) sequentially. The first level, £$\pounds_{1}^{V}$, involves extracting direct features as facial landmarks ($p_{1}$ to $p_{68}$) including eye metrics points ($\varepsilon$) and head position estimation points ($\hbar$). These features are identified using the 68-FLM landmarks detection model provided by the Dlib library. This model has been pre-trained on the i-BUG-300-W dataset where it achieved a (Normalized Root Mean Square Error (NRMSE) of 0.209.[31]
@@
The £$\pounds_{1}^{V}$ also involves extracting numerical features from Eq (1)’s data. This includes the Eye Aspect Ratio ($\xi$) and parameters related to the position of the head $[HP]$, such as pitch $\theta_{x}$, yaw $\theta_{y}$, and roll $\theta_{z}$.
EAR measures the ratio of vertical to horizontal distances between specific eye points. While pitch measures the up-and-down movement of the head, yaw measures the left-to-right rotation of the head, and roll measures the side-to-side tilt of the head. The EAR ($\xi$) is calculated using the facial landmark points of both the left ($\xi_{L}$) and right eyes ($\xi_{R}$), and then their mean is taken.
@@
where $/2]{\sum}$ denotes the computation of the mean and assignment operator. The numerator of these two equations is the distance between the vertical eye landmarks while the denominator is the distance between horizontal eye landmarks. When the eye is open, the EAR is practically constant, but when the eye blinks, it quickly drops to zero. The higher the EAR, the wider the eye is open[32]. For head pose ($HP$) estimation, specific facial landmarks, such as nose tip, chin, corners of the eye, and mouth corners are used. The process assumes no camera distortions during focal length estimation.
The rotation vector $\mathbf{RV}_{[i,j]}^{3\times 1}\big{\}}$, describing the object’s orientation relative to the camera, is computed using the solvePNP function specifically utilizing its iterative approach. The iterative method leverages the Levenberg-Marquardt algorithm[33] [34], a robust optimization technique that utilizes the advantages of both gradient descent and Gauss-Newton methods [35].
The Levenberg-Marquardt algorithm is particularly well-suited for minimizing non-linear least squares problems, such as the reprojection error in pose estimation. Iteratively refining the camera pose parameters ensures convergence towards a local minimum, even in the presence of noise or suboptimal initial estimates. This combination of gradient descent’s stability and the Gauss-Newton method’s rapid convergence enables the iterative approach to achieve both precision and reliability.
The OpenCV first extracts the axis ‘u’ and angle $\theta$ from the rotation vector $\mathbf{RV}_{[i,j]}^{3\times 1}\big{\}}$. It then constructs the skew-symmetric matrix K and calculates the R using the below formula.
@@
The next level, i.e., £$\pounds_{2}^{V}$, involves extracting categorical variables such as eye gaze direction ($\varphi$), head position ($\hbar$), and eye openness category ($\epsilon$) from the numerical variables derived in £$\pounds_{1}^{V}$. Categorical features can capture distinct, non-continuous groups or categories, which often represent real-world concepts more accurately. This distinction can provide meaningful insights that might be lost if the data were treated as numerical. To calculate $\varphi$, the image center ($\odot$) and eye center ($\otimes$) have been used, as follows:
@@
The gaze direction ($\varphi$), computed based on the gaze vector consisting of horizontal (x) and vertical (y) components, indicates whether a person is looking left, right, up, down, or forward based on its values. $\hbar$, computed based on $\theta_{x},\theta_{y}$, and $\theta_{z}$, has 7 classes, including neutral, tilted, or turned in various directions. Eye openness ($\epsilon$), measured by EAR ($\xi$), is categorized into three classes: fully open, partially closed, and closed. All categorical variables depend on thresholds determined experimentally for each variable.
@@ III-B Physiological Feature Extraction Phase
The second phase of feature extraction focuses on extracting the Blood Volume Pulse (BVP) signal from video frames to reflect cardiovascular activity using the Plane-Orthogonal-to-Skin (POS) method[36], a non-invasive technique for rPPG. This method capitalizes on subtle changes in skin color caused by blood flow to estimate the BVP signal.
The £$\pounds_{1}^{P}$ denotes the first level in extracting physiological features, which begins with extracting signals ($S$) by calculating the average RGB values from a specified region of interest (ROI) i.e. $(\Re_{f})$ for the proposed model in each video frame. ROI is identified and extracted using the face_recognition[14] library for accurate localization in each video frame.
@@
where $R_{i}(t),G_{i}(t),$ and $B_{i}(t)$ are the RGB values from the region of interest in the $i^{th}$ video frame, and N is the total number of pixels in the $(\Re_{f})$.
The temporal normalization is then applied to the RGB values by dividing each color signal by its mean over a specific time window. This transformation emphasizes the pulsatile component, primarily captured in the green channel, which is most sensitive to blood flow changes. The normalized signals are calculated as:
$R_{\text{norm}}(t)=\frac{R(t)}{\mu_{R}}\quad G_{\text{norm}}(t)=\frac{G(t)}{\mu_{G}}\quad B_{\text{norm}}(t)=\frac{B(t)}{\mu_{B}}$ (7)
where $\mu_{R}$, $\mu_{G}$, and $\mu_{B}$ are the mean values of the $R(t)$, $G(t)$, and $B(t)$ signals over the time window, respectively. This process reduces noise and variations unrelated to the pulse signal, ensuring the output reflects pulsatile changes due to blood flow.
After normalization, the signal is projected onto a plane orthogonal to the skin tone vector using a transformation matrix. This matrix highlights the pulsatile component, mostly captured in the green channel, sensitive to blood flow changes. The projection is defined as:
$\mathbf{S}_{\text{proj}}(t)=\mathbf{T}\cdot\mathbf{S}_{\text{norm}}(t)$ (8)
$\mathbf{S}_{\text{norm}}(t)=\begin{bmatrix}R_{\text{norm}}(t)\\ G_{\text{norm}}(t)\\ B_{\text{norm}}(t)\end{bmatrix}$ (9)
This transformation emphasizes the pulsatile component, primarily captured in the green channel, which is most sensitive to changes in blood flow.
@@
where $\mathbf{S}_{\text{refined}}(t)$ is the refined signal, $\mathbf{S}_{\text{proj}}(t)$ is the signal’s projection and $\sigma_{S}$ is standard deviation.
@@
After extracting the BVP signal using the POS method, it is further processed to compute meaningful metrics such as heart rate, average peak-to-peak interval, and cumulative sums of systolic and diastolic peaks, providing insights into cardiovascular activity. The RR interval $RR_{i}$ represents the time between consecutive detected heartbeats. The timestamps of the detected peaks are denoted as $t_{i}$ and $t_{i+1}$, where $i$ is the index of the peak. The RR interval for each pair of consecutive peaks is calculated as:
$RR_{i}=t_{i+1}-t_{i}\quad\forall_{i=1}^{N-1}$
where $N$ is the total number of peaks detected in the signal. This gives the interval between each pair of consecutive heartbeats.
The mean RR interval, $\overline{RR}$, is the average of all RR intervals and represents the average time between heartbeats. The formula to calculate the mean RR interval is:
$\overline{RR}=\frac{1}{N-1}\sum_{i=1}^{N-1}RR_{i}$
where $N-1$ represents the number of RR intervals for $N$ detected peaks, and $RR_{i}$ represents the $i$-th RR interval.
The heart rate in beats per minute (BPM) can be derived from the mean RR interval. It is given by:
$\text{BPM}=\frac{60}{\overline{RR}}$
where $\overline{RR}$ is the mean RR interval in seconds, and the factor of 60 converts the interval from seconds to minutes, giving the average heart rate in beats per minute.
The average peak-to-peak interval (PPI), which assesses heart rhythm regularity, is essentially the same as the mean RR interval. It is calculated as:
$\text{PPI}_{\text{avg}}=\frac{1}{N-1}\sum_{i=1}^{N-1}RR_{i}$
Since the PPI is the same as the mean RR interval, we have $\text{PPI}_{\text{avg}}=\overline{RR}$. This value reflects the regularity of the heartbeats during the analysis.
To assess cardiovascular activity, we calculate the cumulative sums of systolic and diastolic peaks. Let $S_{i}$ and $D_{i}$ represent the systolic and diastolic peaks at the $i$-th instance. The cumulative sums of systolic and diastolic peaks are given by:
$CS_{\text{Sys}}=\sum_{i=1}^{N}S_{i}\quad\text{and}\quad CS_{\text{Dia}}=\sum_{i=1}^{N}D_{i}$
where $N$ is the total number of detected systolic and diastolic peaks, and $S_{i}$ and $D_{i}$ represent the values of the systolic and diastolic peaks at each instance.
The cumulative sums provide a measure of the overall cardiovascular activity, with $CS_{\text{Sys}}$ reflecting the sum of systolic events and $CS_{\text{Dia}}$ reflecting the sum of diastolic events during the signal period.